# Planting and MCMC Sampling from the Potts model

### v1

@@ Abstract.
We consider the problem of sampling from the ferromagnetic $q$-state Potts model on the random $d$-regular graph with parameter $\beta>0$. A key difficulty that arises in sampling from the model is the existence of a metastability window $(\beta_{u},\beta_{u}^{\prime})$ where the distribution has two competing modes, the so-called disordered and ordered phases, causing MCMC-based algorithms to be slow mixing from worst-case initialisations.
To this end, Helmuth, Jenssen and Perkins designed a sampling algorithm that works for all $\beta$, when $d\geq 5$ and $q$ large, using cluster expansion methods; more recently, their analysis technique has been adapted to show that random-cluster dynamics mixes fast when initialised more judiciously. However, a well-known bottleneck behind cluster-expansion arguments is that they inherently only work for large $q$, whereas it is widely conjectured that sampling is possible for all $q,d\geq 3$. The only result so far that applies to general $q,d\geq 3$ is by Blanca and Gheissari who showed that the random-cluster dynamics mixes fast for $\beta<\beta_{u}$. For $\beta>\beta_{u}$ however, certain correlation phenomena emerge because of the metastability which have in general been hard to handle, especially for small values of $q$ and $d$.
Our main contribution is to perform a delicate analysis of the Potts distribution and the random-cluster dynamics that goes beyond the threshold $\beta_{u}$. We use planting as the main tool in our proofs, a technique used in the analysis of random CSPs to capture how the space of solutions is correlated with the structure of the random instance. While planting arguments provide only weak sampling guarantees generically, here we instead mesh planting with the analysis of random-cluster dynamics to obtain significantly stronger guarantees. We are thus able to show that the random-cluster dynamics initialised from all-out mixes fast for all integers $q,d\geq 3$ beyond the uniqueness threshold $\beta_{u}$; in fact, our analysis works all the way up to the critical threshold $\beta_{c}\in(\beta_{u},\beta_{u}^{\prime})$ where the dominant mode switches from disordered to ordered. Our arguments also apply to the ordered regime $\beta>\beta_{c}$ where we obtain an algorithm that refines significantly the range of $q,d$.
For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ 1. Introduction
For an integer $q\geq 3$ and a graph $G=(V,E)$, the $q$-state Potts model on $G$ with parameter $\beta$ is a probability distribution $\mu=\mu_{G,q,\beta}$ on $[q]^{V}$ where $[q]=\{1,\ldots,q\}$. Each configuration $\sigma\in[q]^{V}$ has weight $\mu(\sigma)=\mathrm{e}^{\beta m(\sigma)}/Z_{G}$ where $m(\sigma)$ is the number of monochromatic edges under $\sigma$; the normalising factor $Z_{G}$ is the partition function. Throughout, we consider the “ferromagnetic” case $\beta>0$, where configurations with many monochromatic edges are favoured in the distribution.
@@
From a computational complexity perspective, sampling from the ferromagnetic Potts model has various twists relative to other similar models and the complexity of the sampling problem is widely open. On general graphs, the problem is #BIS-hard for any fixed $\beta>0$ [Leslie1]; for graphs of max degree $d$ (where $d$ is a fixed integer), the problem becomes #BIS-hard when $\beta>\beta_{c}$ [BIShard], where $\beta_{c}=\ln\frac{q-2}{(q-1)^{1-2/d}-1}$ is known as the ordered/disordered threshold (we will discuss this in more detail shortly). It is also conjectured that the problem admits a polynomial time algorithm when $\beta<\beta_{c}$ but this is open in general; in fact, many of the standard tools that are used to analyse Markov chains provably fail well below the $\beta_{c}$ threshold. To understand what is going on, the quintessential example is the random $d$-regular graph, which underpins all the relevant phenomena behind this picture. For $dn$ even, we let $\mathbb{G}=\mathbb{G}_{n,d}$ be a graph chosen uniformly at random from the set of all $d$-regular graphs with $n$ vertices. We use “whp” as a shorthand for “with probability $1-o(1)$ as $n$ grows large”.
@@
For a typical random graph $\mathbb{G}$, it is known that a sample from the Potts distribution $\mu_{\mathbb{G}}$ is “disordered” for $\beta<\beta_{c}$, and “ordered” for $\beta>\beta_{c}$; roughly, this means that the colours in the former case appear equally often whereas in the second case one of the colours strictly dominates over the rest (see Lemma 2.1 for the formal statement). The intricate feature of the Potts model on the random $d$-regular graph is that these competing modes are both present as “local maxima” throughout a window $(\beta_{u},\beta_{u}^{\prime})$ containing $\beta_{c}$ even though only one of them has the vast majority of the mass (except at $\beta=\beta_{c}$ itself, where both modes appear with constant probability). This already poses problems to standard MCMC algorithms such as Glauber and Swendsen-Wang dynamics since the simultaneous presence of the modes causes exponentially slow mixing from worst-case initialisations, see [RCM-Helmuth2020, coja2023]. At a more conceptual level, any sampling algorithm has to take into account the presence of the other mode which obliterates standard analysis tools (e.g., correlation decay/spectral independence).
@@
To this end, Helmuth, Jenssen and Perkins [RCM-Helmuth2020] introduced a cluster expansion technique that allowed them to control more precisely how much a typical sample differs from the corresponding mode and obtained an algorithm based on the interpolation method for all $\beta>0$ when $d\geq 5$ and $q\geq d^{\Omega(d)}$; their algorithm applies more generally to expander graphs (under some mild conditions). Following a series of developments [SinclairsGheissari2022, RClattice], this cluster expansion expansion argument has been converted into an MCMC algorithm [RCrandom], using the random-cluster dynamics with appropriate initialisation (see Section 2.2 for details). See also [BlaGhe, Blanca2] for closely related results on the random $d$-regular graph (and lattices) that apply for large $\beta$.
@@
One bottleneck of the cluster expansion technique is that it relies, roughly, on controlling how much a configuration differs from the max-energy configurations and arguing that the difference is relatively small, so it is typically most accurate under appropriate limits (e.g. fixing $d$ and taking $q$ large). By contrast, it is conjectured that sampling on the random regular graph should be possible for all $q,d\geq 3$ (and all $\beta>0$), and hence more precise tools are needed to settle the picture for small $q,d$. The only result so far that works for general $q,d\geq 3$ up to the optimal threshold is by Blanca and Gheissari [BG] who showed that random-cluster dynamics mixes in $O(n\log n)$ time when $\beta<\beta_{u}$, for arbitrary initialisation. Note that for $\beta>\beta_{u}$ the results of [coja2023] give worst-case initialisations that slow down the mixing time to $\mathrm{e}^{\Omega(n)}$; more generally, for $\beta>\beta_{u}$, the presence of the ordered mode imposes correlation phenomena that are hard to handle, especially for small values of $q,d$.
1.1. Results.
@@
Our main contribution is to do a delicate analysis of the Potts distribution and the random-cluster dynamics that goes beyond the threshold $\beta_{u}$, using planting. Planting is a technique used in the analysis of random CSPs to capture how the space of solutions is correlated with the structure of the random instance, see [Plant1, Plant2, Plant3] for some applications. In terms of sampling however, there is no recipe for converting planting methods to efficient sampling algorithms, and typically some extra sampling step is needed on top, see for example [Plant1a, Plant1b] for such approaches; even so, the resulting sample usually has some accuracy limitation, i.e., the TV-distance from the target distribution cannot be made arbitrarily small (typically $n^{-\Omega(1)}$). Here, by meshing planting with the analysis of random-cluster dynamics, we obtain significantly stronger guarantees.
@@
Using this more detailed framework, we give an algorithm to sample from the Potts distribution on the random $d$-regular graph $\mathbb{G}$ when $\beta<\beta_{c}$ for all integers $q,d\geq 3$; as we will explain in more detail later the algorithm is just running random-cluster dynamics for $O(n\log n)$ steps, initialised appropriately (from the “empty” configuration).
@@
Our analysis framework also gives more precise results in the ordered regime $\beta>\beta_{c}$ where we improve upon various restrictions from previous works [UG2, UG1, RCM-Helmuth2020, RCrandom]. For example, [RCM-Helmuth2020, RClattice] applied when $q\geq d^{\Omega(d)}$ (and $d\geq 5$), whereas the recent work by Carlson et al [UG2] (see also [UG1]) applied when $d\geq d_{0}$ and $q\geq d^{c}$ for some large constants $d_{0},c$. Using planting, we obtain a sampling result for all $d\geq 3$ when $q$ is roughly larger than $d^{5}$. In order to improve upon the range of parameters, we essentially require a smaller lower bound on the size of the giant component needed to ensure that so-called polymers are small. We show the following.
@@ 1.2. Proof outline
Our algorithms are based on the random-cluster (RC) representation of the Potts model and running the analogue of Glauber dynamics in that setting. So, in the next Preliminaries section we define the RC model and the relevant random-cluster chain. We also give a more detailed overview of the phases of the Potts model, and how planting gives a handle for the disordered and ordered settings. Most of these parts are largely imported from [coja2023, BIShard].
In Section 3, we convert the planting for Potts into suitable planting for the RC model, giving the analogue of the ordered and disordered phases in that setting. Then, we state the main contribution of this work which is establishing weak spatial mixing within the disordered and ordered phases (a notion that originated in [SinclairsGheissari2022]). Sections 4 and 5 give the technical core of our arguments.
@@
Roughly, in the disordered regime, the key ingredient is to show that the monochromatic edges do not typically form large components. For $\beta<\beta_{u}$, much of this proof strategy was present in the work of Blanca and Gheissari [BG] where they used a very careful random-graph revealing process coupled with the evolution of random-cluster dynamics. Using planting, we simplify significantly this step since planting allows us to capture the correlation between the random graph and a sample from stationarity more accurately, reducing the study of the component structure to showing that the planted random graph is subcritical. Having this in place (Lemma 4.2), the rest of the proof is more streamlined after some small reworking of certain correlation decay estimates up to the ordered/disordered threshold $\beta_{c}$ (Lemma 4.3).
@@
In the ordered regime, the planted model is instead supercritical which overall complicates the arguments. The key ingredient there is to show that every vertex $v$ is surrounded by close-by vertices (at distance $\ell<\tfrac{1}{2}\log_{d-1}n$) that belong to the “giant” component (a so-called wired boundary). Planting here allows us to do a head-on analysis of the probability that, in a typical sample from stationarity, there exists a path of length $\ell$ that does not include any vertex in the giant. We get a tight bound for this, so then it is a matter of taking $q$ large ($q\geq(5d)^{5}$) so that a union bound over the paths starting at $v$ gives the existence of the desired wired boundary (in a typical configuration).
1.3. Further Discussion.
@@
We remark that the planting arguments that we use in this paper follow by careful first and second moment considerations in [BIShard], which have been carried out for the ferromagnetic Potts model. A somewhat similar idea based on moment analysis can be found in [Chen] for sampling independent sets and proper colourings in random bipartite graphs (though there the approach ultimately relies on cluster-expansion type of arguments [BIS1, BIS2]). Our more direct planting approach can likely be carried out in the independent setting as well and obtain sharper bounds using Glauber dynamics, but as it is pointed out in [Chen, Lemma 10] it seems unlikely that it will achieve better asymptotic estimates in terms of the degree $d$.
@@
It is also relevant to note that the results we stated earlier for the cluster-expansion technique and fast mixing of the random-cluster dynamics in the literature apply to non-integer $q$, in the random-cluster representation. It is conceivable that a similar planting to the one we consider here for the Potts model can be carried out for non-integer $q$ in the random-cluster representation, though the related first and second moment computations will likely require careful arguments for the distribution of the number of components in random graph percolation; see [bencs2022random] for some related developments.
@@ 2.1. The configuration model
To obtain a handle on the random $d$-regular $\mathbb{G}=\mathbb{G}_{n,d}$, we work throughout in the configuration model. Formally, for integers $n,d\geq 1$ with $dn$ even, let $[n]\times[d]$ be a set of “half-edges”; we denote by $G\sim\mathcal{G}_{n,d}$ a (multi)graph on vertex set $[n]$ whose edges are obtained by taking a uniformly random perfect matching of the half-edges; we add an edge between $u,v\in[n]$ in $G$ whenever the half-edges $(u,i)$ and $(v,j)$ are matched together for some $i,j\in[d]$. For brevity, instead of $G\sim\mathcal{G}_{n,d}$ we sometimes use $\boldsymbol{G}$ to denote the random (multi)graph chosen. It is a standard fact that any property that holds whp over $\boldsymbol{G}$ also holds whp over $\mathbb{G}$ (see, e.g., [JLR]), so henceforth we focus on the configuration model $\boldsymbol{G}$.

### v2

@@ Abstract.
We consider the problem of sampling from the ferromagnetic $q$-state Potts model on the random $d$-regular graph with parameter $\beta>0$. A key difficulty that arises in sampling from the model is the existence of a “metastability” window $\beta\in(\beta_{u},\beta_{u}^{\prime})$, where roughly the distribution has two competing modes, the so-called disordered and ordered phases. This causes classical Markov-chain algorithms to be slow mixing from worst-case initialisations. Nevertheless, Helmuth, Jenssen and Perkins (SODA ’19) designed a sampling algorithm that works for all $\beta$, when $d\geq 5$ and $q=d^{\Omega(d)}$, using polymers and cluster expansion methods; more recently, their analysis technique has been adapted to show that a Markov chain (random-cluster dynamics) mixes fast when initialised appropriately, in the same regime of $q,d,\beta$.
Despite these positive algorithmic results, a well-known bottleneck behind cluster-expansion arguments is that they inherently only work for large $q$, whereas it is widely conjectured that sampling on the random $d$-regular graph is possible for all $q,d\geq 3$. The only result so far that applies to general $q,d\geq 3$ is by Blanca and Gheissari who showed that the random-cluster dynamics mixes fast in the “uniqueness” regime $\beta<\beta_{u}$ where roughly only the disordered mode exists. For $\beta\geq\beta_{u}$ however, a second subdominant mode emerges creating bottlenecks and giving rise to correlations which have been hard to handle, especially for small values of $q$ and $d$.
Our main contribution is to perform a delicate analysis of the Potts distribution and the random-cluster dynamics that goes beyond the threshold $\beta_{u}$. We use planting as the main tool, a technique used in the analysis of random CSPs to capture how the space of solutions is correlated with the structure of the random instance. While planting arguments provide only weak sampling guarantees generically, here we instead combine planting with the analysis of random-cluster dynamics to obtain significantly stronger guarantees. We are thus able to show that the random-cluster dynamics initialised from all-out mixes fast for all integers $q,d\geq 3$ beyond the uniqueness threshold $\beta_{u}$, all the way to the optimal threshold $\beta_{c}\in(\beta_{u},\beta_{u}^{\prime})$ where the dominant mode switches from disordered to ordered. A more involved analysis also applies to the ordered regime $\beta>\beta_{c}$ where we obtain an algorithm for all $d\geq 3$ and $q\geq(5d)^{5}$, improving significantly upon the previous range of $q,d$ by Carlson, Davies, Fraiman, Kolla, Potukuchi, and Yap (FOCS’22).
For the purpose of Open Access, the author has applied a $C$ BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
@@ 1. Introduction
For an integer $q\geq 3$ and a graph $G=(V,E)$, the $q$-state Potts model on $G$ with parameter $\beta$ is a probability distribution $\mu=\mu_{G,q,\beta}$ on $q]^{V}$ where $q]=\{1,\ldots,q\}$. Each configuration $\sigma\in[q]^{V}$ has weight $\mu(\sigma)=\mathrm{e}^{\beta m(\sigma)}/Z_{G}$ where $m(\sigma)$ is the number of monochromatic edges under $\sigma$; the normalising factor $Z_{G}$ is the partition function. Throughout, we consider the “ferromagnetic” case $\beta>0$, where configurations with many monochromatic edges are favoured in the distribution.
@@
From a computational complexity perspective, sampling from the ferromagnetic Potts model has various twists relative to other similar models and the complexity of the sampling problem is widely open. On general graphs, the problem is #BIS-hard for any fixed $\beta>0$ [23]; for graphs of max degree $d$ (where $d$ is a fixed integer), the problem becomes #BIS-hard when $\beta>\beta_{c}$ [19], where $\beta_{c}=\ln\frac{q-2}{(q-1)^{1-2/d}-1}$ is known as the ordered/disordered threshold (we will discuss this in more detail shortly). It is also conjectured that the problem admits a polynomial time algorithm when $\beta<\beta_{c}$ but this is open in general; in fact, many of the standard tools that are used to analyse Markov chains provably fail well below the $\beta_{c}$ threshold. The quintessential example here is the random $d$-regular graph, which underpins all the relevant phenomena behind this picture. For $dn$ even, we let $\mathbb{G}=\mathbb{G}_{n,d}$ be a graph chosen uniformly at random from the set of all $d$-regular graphs with $n$ vertices. We use “whp” as a shorthand for “with probability 1o(1)1-o(1) as $n$ grows large”.
@@
For a typical random graph $\mathbb{G}$, it is known that a sample from the Potts distribution $\mu_{\mathbb{G}}$ is “disordered” for $\beta<\beta_{c}$, and “ordered” for $\beta>\beta_{c}$; roughly, this means that the colours in the former case appear equally often whereas in the second case one of the colours strictly dominates over the rest (see Lemma 2.1 for the formal statement). The intricate feature of the Potts model on the random $d$-regular graph is that these competing modes are both present as “local maxima” throughout a window $(\beta_{u},\beta_{u}^{\prime})$ containing $\beta_{c}$ even though only one of them has the vast majority of the mass (except at $\beta=\beta_{c}$ itself, where both modes appear with constant probability). This already poses problems to standard MCMC algorithms such as Glauber and Swendsen-Wang dynamics since the simultaneous presence of the modes causes exponentially slow mixing from worst-case initialisations, see [25, 14]. At a more conceptual level, any sampling algorithm has to take into account the presence of the other mode which obliterates from the very start standard analysis tools (e.g., correlation decay/spectral independence).
@@
Helmuth, Jenssen and Perkins [25] introduced a cluster-expansion technique that allowed them to control more precisely how much a typical sample differs from the corresponding mode, and obtained an algorithm based on the interpolation method for all $\beta>0$ when $d\geq 5$ and $q\geq d^{\Omega(d)}$; their algorithm applies more generally to expander graphs (under some mild conditions). See also the works [16, 11, 12] for various refinements of their method. Following a series of developments [20, 21], this cluster expansion expansion argument has been converted into an MCMC algorithm (for large $d$ and $q$) [18] using the random-cluster dynamics with appropriate initialisation that avoids the bottlenecks in the distribution (see Section 2.2 for details). See also [10, 8, 22] for closely related results on the random $d$-regular graph (and lattices) that apply for large $\beta$.
@@
Despite these positive algorithmic developments, the cluster expansion technique that underpins the analysis of these algorithms is inherently “perturbative”, relying roughly on controlling how much a typical configuration differs from the max-energy configurations and arguing that the difference is relatively small. In the case of the Potts model, the resulting estimates are accurate only when $q$ is quite large relative to $d$. By contrast, it is conjectured that sampling on the random regular graph should be possible for all $q,d\geq 3$ (and all $\beta>0$), and hence more precise tools are needed to settle the picture for small $q,d$. The only result so far that works for general $q,d\geq 3$ is by Blanca and Gheissari [9] who showed that random-cluster dynamics mixes in $O(n\log n)$ time when $\beta<\beta_{u}$, for arbitrary initialisation. Note that for $\beta>\beta_{u}$ the results of [14] give worst-case initialisations that slow down the mixing time of the Glauber dynamics to $\mathrm{e}^{\Omega(n)}$; more generally, for $\beta>\beta_{u}$, the presence of the ordered mode imposes correlation phenomena that are hard to handle, especially for small values of $q,d$.
1.1. Results
@@
Our main contribution is to do a delicate analysis of the Potts distribution and the random-cluster dynamics that goes beyond the threshold $\beta_{u}$, using planting. Planting is a technique used in the analysis of random CSPs to capture how the space of solutions is correlated with the structure of the random instance, see [1, 15, 5] for some applications. In terms of sampling however, there is no recipe for converting planting methods to efficient sampling algorithms, and typically some extra sampling step is needed on top, see for example [2, 17] for such approaches; even so, the resulting sample usually has some accuracy limitation, i.e., the TV-distance from the target distribution cannot be made arbitrarily small (typically it is required to be at least an inverse polynomial in $n$). Here, we instead use planting as a means to obtain sharp estimates that are used in the analysis of random-cluster dynamics; we thus show fast convergence of the latter, yielding automatically strong sampling guarantees. A different sampling idea that uses planting in a somewhat related manner has recently been considered in [26], in the context of sampling from spherical spin glasses.
@@
Using this planting framework, we give an algorithm to sample from the Potts distribution on the random $d$-regular graph $\mathbb{G}$ that covers the optimal range of “high temperatures” $\beta<\beta_{c}$ for all integers $q,d\geq 3$; as we will explain in more detail later the algorithm is running random-cluster dynamics for $O(n\log n)$ steps, initialised appropriately (from the “empty” configuration).
@@
Our analysis framework also applies to the ordered regime $\beta>\beta_{c}$ where we improve significantly upon various restrictions from previous works [11, 16, 25, 18]. For example, [25, 21] applied when $q\geq d^{\Omega(d)}$ (and $d\geq 5$), whereas the recent work by Carlson et al [11] (see also [16]) applied when $d\geq d_{0}$ and $q\geq d^{c}$ for some large constants $d_{0},c$. Using planting, we obtain a sampling result for all $d\geq 3$ when $q$ is roughly larger than $d^{5}$. We show the following.
@@ 1.2. Proof outline
Our algorithms are based on a well-known edge representation of the Potts model, the random-cluster (RC) model, where configurations are edge subsets weighted by the number of edges present and the number of induced components. We run the random-cluster dynamics (the analogue of Glauber dynamics in this edge setting), starting from the empty set of edges in the disordered regime $\beta<\beta_{c}$ and the full set of edges in the ordered regime $\beta>\beta_{c}$. The reason for working in the RC representation is that, in contrast to the Potts model, the RC model enjoys certain monotonicity properties that facilitate tracking the evolution of Glauber dynamics. In the next Preliminaries section we define the RC model and the relevant random-cluster chain. We also give a more detailed overview of the phases of the Potts model, and how planting gives a handle for the disordered and ordered settings. Most of these parts are largely imported from [14, 19]. In Section 3, we convert the planting for Potts into suitable planting for the RC model, giving the analogue of the ordered and disordered phases in that setting. Then, we state the main contribution of this work which is establishing weak spatial mixing within the disordered and ordered phases (a notion that originated in [20]). Sections 4 and 5 give the technical core of our arguments.
@@
Roughly, in the disordered regime (Section  4), the key ingredient is to show that the dynamics converges to a set of configurations that do not form large components. For $\beta<\beta_{u}$, Blanca and Gheissari [9] accomplished this by a very careful random-graph revealing process coupled with the evolution of random-cluster dynamics. Using planting, we are able to capture precisely the correlation between the random graph and a sample from stationarity for $\beta<\beta_{c}$, reducing the study of the component structure to showing that the planted random graph is subcritical. Using the monotonicity of the RC model, this implies that the RC dynamics initialised from the empty set (“all-out”) enjoys the same component structure as the stationary distribution (for exponentially many steps). Having this in place (Lemma 4.2), the rest of the proof is more streamlined after some small reworking of certain correlation decay estimates up to the ordered/disordered threshold $\beta_{c}$ (Lemma 4.3).
@@
In the ordered regime (Section 5), the planted model is instead supercritical and a typical configuration has a “giant” component, whose presence complicates the arguments significantly. The goal now is roughly to show that bichromatic edges do not typically form large components; more precisely, the key ingredient is to show that every vertex $v$ is surrounded by close-by vertices (each at some distance $\ell<\tfrac{1}{2}\log_{d-1}n$) that belong to the giant component, a so-called “wired” boundary. Planting here allows us to do a head-on analysis of the probability that, in a typical sample from stationarity, there exists a path of length $\ell$ that does not include any vertex in the giant. We get a tight bound for this using careful first moment arguments and percolation bounds; then, for $q\geq(5d)^{5}$, a union bound over the paths starting at $v$ gives the existence of the desired wired boundary (in a typical configuration). Relative to the earlier works [25, 16, 11, 18], bichromatic components roughly correspond to so-called polymers; our techniques show that a smaller lower bound on the size of the giant component suffices in order to ensure that the polymers are not large.
1.3. Further Discussion
@@
The planting arguments that we use in this paper follow by careful first and second moment considerations in [19], which have been carried out for the ferromagnetic Potts model, see also [4] for similar-flavored “silent” planting results. The moment analysis has previously been used in [13] for sampling independent sets and proper colourings in random bipartite graphs, though there the approach ultimately relies on cluster-expansion type of arguments [28, 12]. Our more direct planting approach can be carried out for sampling independent sets on random bipartite graphs (that shares similar monotonicity and planting properties) and obtain sharper bounds for sampling using Glauber dynamics. Note however that, as pointed out in [13, Lemma 10], it is unlikely that current polymer-based techniques will achieve better asymptotic estimates in terms of the degree $d$. Our technique should also apply to the $G(n,d/n)$ setting (for arbitrary $q,d>1$) provided that the analogous “silent” planting is in place; this seems more easily feasible in the disordered regime but more difficult for the ordered regime where the Poisson-tree neighbourhoods are expected to cause bigger variance.
@@
It is also relevant to note that the results we stated earlier for the cluster-expansion technique and fast mixing of the random-cluster dynamics in the literature apply to non-integer $q$, in the random-cluster representation. It is conceivable that a similar planting to the one we consider here for the Potts model can be carried out for non-integer $q$ in the random-cluster representation, though the related first and second moment computations will likely require careful arguments for the distribution of the number of components in random graph percolation; see [6] for some related developments.
@@ 2.1. The configuration model
To obtain a handle on the random $d$-regular $\mathbb{G}=\mathbb{G}_{n,d}$, we work throughout in the configuration model. Formally, for integers $n,d\geq 1$ with $dn$ even, let $n]\times[d]$ be a set of “half-edges”; we denote by $G\sim\mathcal{G}_{n,d}$ a (multi)graph on vertex set $[n]$ whose edges are obtained by taking a uniformly random perfect matching of the half-edges; we add an edge between $u,v\in[n]$ in $G$ whenever the half-edges $(u,i)$ and $(v,j)$ are matched together for some $i,j\in[d]$. For brevity, instead of $G\sim\mathcal{G}_{n,d}$ we sometimes use $\boldsymbol{G}$ to denote the random (multi)graph chosen. Any property that holds whp over $\boldsymbol{G}$ also holds whp over $\mathbb{G}$ (see, e.g., [27]), so henceforth we focus on the configuration model $\boldsymbol{G}$.